# The Study of the Canonical forms of Killing tensor in vacuum with $\Lambda$

### v1

@@ 4.2.2 The diagonalized form of the Canonical forms
${K^{0\mu}}_{\nu}=\begin{cases}\lambda_{1}=-(\lambda_{2}+\lambda_{7})\ ;\ \begin{pmatrix}\lambda_{1}&1&0&0\\ 0&\lambda_{1}&1&0\\ 0&0&\lambda_{1}&0\\ 0&0&0&-(\lambda_{2}-\lambda_{7})\end{pmatrix}\\ \lambda_{1}=-(\lambda_{2}-\lambda_{7})\ ;\ \begin{pmatrix}\lambda_{1}&1&0&0\\ 0&\lambda_{1}&1&0\\ 0&0&\lambda_{1}&0\\ 0&0&0&-(\lambda_{2}+\lambda_{7})\end{pmatrix}\end{cases}$ (Case 0)
${K^{1\mu}}_{\nu}=\begin{pmatrix}\lambda_{1}&1&0&0\\ 0&\lambda_{1}&0&0\\ 0&0&-(\lambda_{2}+\lambda_{7})&0\\ 0&0&0&-(\lambda_{2}-\lambda_{7})\end{pmatrix}$ (Case 1)
${K^{2\mu}}_{\nu}=\begin{pmatrix}\lambda_{0}+\lambda_{1}&0&0\\ 0&\lambda_{0}-\lambda_{1}&0&0\\ 0&0&-(\lambda_{2}+\lambda_{7})&0\\ 0&0&0&-(\lambda_{2}-\lambda_{7})\end{pmatrix}$ (Case 2)
${K{{}^{3\mu}}}_{\nu}=\begin{pmatrix}\lambda_{0}+i\lambda_{1}&0&0\\ 0&\lambda_{0}-i\lambda_{1}&0&0\\ 0&0&-(\lambda_{2}+\lambda_{7})&0\\ 0&0&0&-(\lambda_{2}-\lambda_{7})\end{pmatrix}$ (Case 3)
@@ 5 Solutions of $K^{0}_{\mu\nu}$ form
$K^{0}_{\mu\nu}=\lambda_{1}(l_{\mu}n_{\nu}+n_{\mu}l_{\nu})-p(n_{\mu}m_{\nu}+m_{\mu}n_{\nu})-\bar{p}(n_{\mu}\bar{m}_{\nu}+\bar{m}_{\mu}n_{\nu})\\ +\lambda_{7}(m_{\mu}m_{\nu}+\bar{m}_{\mu}\bar{m}_{\nu})+\lambda_{2}(m_{\mu}\bar{m}_{\nu}+\bar{m}_{\mu}m_{\nu});\hskip 5.69046ptp=-\bar{p}=\pm 1$ (35)
@@
We are already aware that $Q+1=0$ or $Q-1=0$. These two possibilities are the necessary conditions to the existence of a 3-dimensional Jordan block and also guarantee the existence of one timelike and three spacelike eigenvectors. Let us proceed with the first choice where $Q+1=0$. Then, the equations regardless the value of the factor $p$ yield
@@ 5.2 A unique Petrov type D solution
Finally, we performed the exact same derivations for $Q-1=\lambda_{1}+\lambda_{2}-\lambda_{7}=0$, and we arrived at the same Petrov type D, with the same characteristics of the solution to be hold.
@@
A unique Petrov type D solution with a non-geodesic, shearing and diverging null congruence $n^{\mu}$ admits $K^{0}_{\mu\nu}$ canonical form of Killing tensor.
@@ 6 Solutions of $K^{1}_{\mu\nu}$ form
With these simplifications, we obtain a Jordan form of Killing tensor, as opposed to a diagonalized form with two double eigenvalues. Besides, it is widely known that the Jordan canonical form of a matrix embodies all the similar matrices of the family of matrices with the same eigenvalues except the “unique” member of this family, the diagonalized member of the family [Strang, 2022]. We shall proceed now with the Killing equation serves as a starting point,
@@ 6.2.3 Class III: $\mu\neq 0=\tau$
Petrov type N solution with a shear-free (non-diverging in one case) null congruence admits $K^{1}_{\mu\nu}$ canonical form of Killing tensor with $\lambda_{7}=0$.
@@ 7.2.3 Class III: $\tau=0$
Petrov types III, N solutions with a shear-free, diverging (and in some cases geodesic) null congruence admit both $K^{2}_{\mu\nu},K^{3}_{\mu\nu}$ canonical forms of Killing tensor with $\lambda_{7}=0$.
@@
Petrov type D solution with a shear-free, diverging and non-geodesic null congruence admit both $K^{2}_{\mu\nu},K^{3}_{\mu\nu}$ canonical forms of Killing tensor with $\lambda_{7}=0$.
@@ 8.2.3 Class III: $\rho=0\neq\pi$
Petrov type N and D solutions with a geodesic, shearing and diverging null congruence admit $K^{4}_{\mu\nu}$ canonical form of Killing tensor with $\lambda_{0}=0$.
@@
Petrov type N and III solutions with a geodesic, non-shearing and non-diverging null congruence admit $K^{4}_{\mu\nu}$ canonical form of Killing tensor with $\lambda_{0}=0$.
@@ 9 Discussion and Conclusions
The Study of the Canonical forms of Killing tensor is initiated by considering the possibility to obtain more interesting spacetimes using the canonical forms as an initial premise. In order to prove this statement properly we should rather have extracted the corresponding solutions that admit the canonical forms of Killing tensor in electro-vacuum, where the Einstein-Maxwell tensor $F_{\mu\nu}$ would be present and comparing to those of the work of Hauser-Malhiot [Hauser and Malhiot, 1976]. This work will be conducted in the immediate future. However, quite useful conclusions were extracted during this investigation (Vacuum with $\Lambda$) concerning this hypothesis.
@@
The latter choice falls into the concept of the symmetric null tetrads that was manifested in [Debever and McLenaghan, 1981]. Using the free rotation parameters $a$ and $b$ Debever managed to obtain the following simplifications that led him to the most general type D solution with a geodesic, shear-free null congruence [Debever et al., 1984] assuming that the PND of Einstein-Maxwell tensor aligned to the PND of Weyl tensor. Namely, the following relations determine the rotation parameters $a$ and $b$
@@
The interesting part arises when we derived the Petrov types of the extracted solutions assuming the existence of the Killing tensor of our Paradigm both ways. Firstly we implied the symmetric null tetrad concept, afterwards the null rotation methodology of Chapter 3. With fundamental derivations it is really easy to prove that using the symmetric null tetrad concept one extracts a unique type D solution in vacuum with $\Lambda$. In fact, this type D solution is a limitation in vacuum with $\Lambda$ of Debever’s solution [Debever et al., 1984] and Papakostas’ solution [Papakostas, 1998], wherein both solutions were obtained based on the aforementioned concept.
Although, applying the methodology of Chapter 3 we result in two conformally flat solutions  [^8] The proof of this statement is not given here, although it is easy for the reader to verify this using the IC of the Paradigm (relations (11) in [Papakostas, 1998])., wherein the implication of the null rotation just annihilates the parameter $p$. Basically this outcome shows that we obtain more general Petrov types solutions (type D) abolishing the arbitrariness of the tetrad. Subsequently, a question arises, which is the most general transformation that we must apply in order to acquire the most general Petrov type solution?
Moving forward, another important matter was brought to surface by this study, is that the canonical forms appear to be more fruitful instead of the Killing tensor of our Paradigm satisfying the conjecture of Hauser-Malhiot [Hauser and Malhiot, 1976]. This is evident by the vast variety of the different Petrov types were obtained assuming the $K^{2},K^{3}$ Killing forms and $K^{4}$ form by extension  [^9] Once more, we denote that $K^{4}$ form is not a typical canonical form but it is originated by the reduction of $K^{1},K^{2},K^{3}$ where $\lambda_{0}=0$.. Also, in reference to the work of Van den Bergh [Van den Bergh, 2017] our type D solutions, one with a shear-free and non-geodesic null congruence (Theorem 4) and the other with a shearing and geodesic null congruence (Theorem 5) might be new type D solutions. In fact, they could be possible reductions (in vacuum with $\Lambda$) of the undiscovered solutions with PND of Einstein-Maxwell tensor aligned (or non-aligned) with the PND of Weyl tensor.

### v2

@@ 4.2.4 The diagonalized form of the Canonical forms
${K^{0}_{\mu}}^{\nu}=\begin{pmatrix}\lambda_{1}&1&0&0\\ 0&\lambda_{1}&1&0\\ 0&0&\lambda_{1}&0\\ 0&0&0&-(\lambda_{2}+\lambda_{7})\end{pmatrix};\hskip 2.84544pt\lambda_{1}=-(\lambda_{2}-\lambda_{7})$ (Case 0)
${K^{1}_{\mu}}^{\nu}=\begin{pmatrix}\lambda_{1}&1&0&0\\ 0&\lambda_{1}&0&0\\ 0&0&-(\lambda_{2}+\lambda_{7})&0\\ 0&0&0&-(\lambda_{2}-\lambda_{7})\end{pmatrix}$ (Case 1)
${K^{2}_{\mu}}^{\nu}=\begin{pmatrix}\lambda_{1}+\lambda_{0}&0&0\\ 0&\lambda_{1}-\lambda_{0}&0&0\\ 0&0&-(\lambda_{2}+\lambda_{7})&0\\ 0&0&0&-(\lambda_{2}-\lambda_{7})\end{pmatrix}$ (Case 2)
${K^{3}_{\mu}}^{\nu}=\begin{pmatrix}\lambda_{1}+i\lambda_{0}&0&0\\ 0&\lambda_{1}-i\lambda_{0}&0&0\\ 0&0&-(\lambda_{2}+\lambda_{7})&0\\ 0&0&0&-(\lambda_{2}-\lambda_{7})\end{pmatrix}$ (Case 3)
@@ 5 Solutions of $K^{0}_{\mu\nu}$ form
$K^{0}_{\mu\nu}=\lambda_{1}(l_{\mu}n_{\nu}+n_{\mu}l_{\nu})-p(n_{\mu}m_{\nu}+m_{\mu}n_{\nu})-\bar{p}(n_{\mu}\bar{m}_{\nu}+\bar{m}_{\mu}n_{\nu})\\ +\lambda_{7}(m_{\mu}m_{\nu}+\bar{m}_{\mu}\bar{m}_{\nu})+\lambda_{2}(m_{\mu}\bar{m}_{\nu}+\bar{m}_{\mu}m_{\nu});\hskip 5.69046ptp=-\bar{p}=\pm i$ (43)
@@
We are already aware that $Q+1=0$. Then, the equations regardless the value of the factor $p$ yield
@@ 5.2 A unique Petrov type D solution
Assuming the existence of $K^{0}_{\mu\nu}$ canonical form of Killing tensor in vacuum with $\Lambda$ a Petrov type D solution with a non-geodesic, shearing and diverging null congruence $n^{\mu}$ emerged.
@@ 6 Solutions of $K^{1}_{\mu\nu}$ form
With these simplifications, we obtain a Jordan form of Killing tensor, as opposed to a diagonalized form with two double eigenvalues. Besides, it is widely known that the Jordan canonical form of a matrix embodies all the similar matrices of the family of matrices with the same eigenvalues except the “unique” member of this family, the diagonalized member of the family [34]. We shall proceed now with the Killing equation serves as a starting point,
@@ 6.2.3 Class III: $\mu\neq 0=\tau$
Assuming $K^{1}_{\mu\nu}$ canonical form of Killing tensor with $\lambda_{7}=0$ in vacuum with $\Lambda$ and applying a null rotation around the null tetrad frame Petrov type N solutions with a shear-free (non-diverging in one case) null congruence emerged.
@@ 7.2.3 Class III: $\tau=0$
Assuming $K^{2}_{\mu\nu},K^{3}_{\mu\nu}$ canonical forms of Killing tensor with $\lambda_{7}=0$ in vaccum with $\Lambda$ and applying a null rotation around the null tetrad frame Petrov types III, N solutions with a shear-free, diverging (and in some cases geodesic) null congruence emerged.
@@
Assuming $K^{2}_{\mu\nu},K^{3}_{\mu\nu}$ canonical forms of Killing tensor with $\lambda_{7}=0$ in vacuum with $\Lambda$ and applying a null rotation around the null tetrad frame a unique Petrov type D solution with a shear-free, diverging and non-geodesic null congruence emerged.
@@ 8.2.3 Class III: $\rho=0\neq\pi$
Assuming $K^{4}_{\mu\nu}$ canonical form of Killing tensor with $\lambda_{0}=0$ in vacuum with $\Lambda$ and applying a null rotation around the null tetrad frame Petrov type N and D solutions with a geodesic, shearing and diverging null congruence emerged.
@@
Assuming $K^{4}_{\mu\nu}$ canonical form of Killing tensor with $\lambda_{0}=0$ in vacuum with $\Lambda$ and applying a null rotation around the null tetrad frame Petrov type N and III solutions with a geodesic, non-shearing and non-diverging null congruence emerged.
@@ 9 Discussion and Conclusions
The Study of the Canonical forms of Killing tensor is initiated by considering the possibility to obtain more interesting spacetimes with hidden symmetries using the canonical forms as an initial premise. In order to prove this statement properly we should rather have extracted the corresponding solutions that admit the canonical forms of Killing tensor in electro-vacuum, where the Einstein-Maxwell tensor $F_{\mu\nu}$ would be present and comparing to those of the work of Hauser-Malhiot [6]. This work will be conducted in the immediate future. However, quite useful conclusions were extracted during this investigation (Vacuum with $\Lambda$) concerning this hypothesis.
@@
The latter choice falls into the concept of the symmetric null tetrads that was manifested in [35]. Using the free rotation parameters $a$ and $b$ Debever managed to obtain the following simplifications that led him to the most general type D solution with a geodesic, shear-free null congruence [36] assuming that the PND of Einstein-Maxwell tensor aligned to the PND of Weyl tensor. Namely, the following relations determine the rotation parameters $a$ and $b$
@@
$e=\pm 1$
The interesting part arises when we derived the Petrov types of the extracted solutions assuming the existence of the Killing tensor of our Paradigm both ways. Firstly we implied the symmetric null tetrad concept, afterwards the null rotation methodology of Chapter 3. With fundamental derivations it is really easy to prove that using the symmetric null tetrad concept one extracts a unique type D solution in vacuum with $\Lambda$. In fact, this type D solution is a limitation in vacuum with $\Lambda$ of Debever’s solution [36] and Papakostas’ solution [8], wherein both solutions were obtained based on the aforementioned concept.
Although, applying the methodology of Chapter 3 we result in two conformally flat solutions  [^10] The proof of this statement is not given here, although it is easy for the reader to verify this using the IC of the Paradigm (relations (11) in [8])., wherein the implication of the null rotation just annihilates the parameter $p$. Basically this outcome shows that we obtain more general Petrov types solutions (type D) abolishing the arbitrariness of the tetrad. Subsequently, a question arises, which is the most general transformation that we must apply in order to acquire the most general Petrov type solution?
Moving forward, another important matter was brought to surface by this study, is that the canonical forms appear to be more fruitful instead of the Killing tensor of our Paradigm satisfying the conjecture of Hauser-Malhiot [6]. This is evident by the vast variety of the different Petrov types were obtained assuming the $K^{2},K^{3}$ Killing forms and $K^{4}$ form by extension  [^11] Once more, we denote that $K^{4}$ form is not a typical canonical form but it is originated by the reduction of $K^{1},K^{2},K^{3}$ where $\lambda_{0}=0$.. Also, in reference to the work of Van den Bergh [5] our type D solutions, one with a shear-free and non-geodesic null congruence (Theorem 4) and the other with a shearing and geodesic null congruence (Theorem 5) might be new type D solutions. In fact, they could be possible reductions (in vacuum with $\Lambda$) of the undiscovered solutions with PND of Einstein-Maxwell tensor aligned (or non-aligned) with the PND of Weyl tensor.